# Heat Flow in Semi-Flat Elliptic Fibrations

### v1

@@
Abstract
Motivated by the SYZ picture for the collapsing of elliptic K3 surfaces, we study heat kernels under semi-flat collapse of Ricci-flat Kähler manifolds $(X_{t},g(t))$ fibred by flat $2$-tori $\Pi_{t}:X_{t}\to B$ over a surface $B$ with a finite discriminant set $D$. At the regular locus $B_{\mathrm{reg}}:=B\setminus D$ we assume an exponentially accurate semi-flat product approximation and a uniform vertical spectral gap at the collapse scale. Using the associated normalized lift and fiber-average maps $(I_{t},P_{t})$, we show that for each fixed time $\tau>0$ the fiber-compressed heat operators
$P_{t}e^{-\tau\Delta_{g(t)}}I_{t}$
converge strongly on $L^{2}$ to the heat semigroup $e^{-\tau\Delta_{g_{B}}}$ on the base; equivalently, bilinear pairings of the total-space heat kernel against fiber-constant lifts of test functions converge on any precompact subset of $B_{\mathrm{reg}}$.
The main additional issue is the discriminant: even when $X_{t}$ is smooth, the limiting base geometry is conic near $D$, so test functions that meet $D$ cannot be handled by a purely interior argument on $B_{\mathrm{reg}}$. In wedge charts around $D$ we introduce a canonical conic-renormalized bilinear functional $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ by patching the base heat-kernel pairing away from $D$ with the model heat kernel of the corresponding flat cone near each discriminant point. A conic parametrix shows that $K_{B}^{\mathrm{ren}}$ is well-defined (cutoff- and chart-independent) and that the total-space bilinear pairings converge to $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ for arbitrary $\Phi,\Psi\in C_{c}^{\infty}(B)$ and every fixed $\tau>0$. When $\mathrm{supp}\,\Phi\cup\mathrm{supp}\,\Psi\subset B_{\mathrm{reg}}$, the renormalization is trivial and $K_{B}^{\mathrm{ren}}$ reduces to the usual base heat-kernel pairing. Under exponentially small geometric errors we obtain an explicit exponential convergence rate, uniform for $\tau$ in compact subintervals of $(0,\infty)$. The results extend to Neumann and nonnegatively weighted Robin boundary conditions and globalize by exhaustion of $B_{\mathrm{reg}}$.
1 Introduction and statement of results
1.1 Motivation: heat flow under SYZ-type collapse
@@
1.2 Geometric setting and analytic identifications
@@
1.3 Bilinear pairings and compressed semigroups
@@
$\iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ \langle\Phi,\ (P_{t}T_{t}(\tau)I_{t})\Psi\rangle_{L^{2}(B_{0})}.$ (1)
@@
1.4 Discriminant regime and intrinsic conic renormalization
@@
$K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[\iint_{B\times B}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)d\mu_{g_{B}}(b^{\prime})+\sum_{j=1}^{N}\iint_{U_{j}\times U_{j}}(\chi_{j}\otimes\chi_{j})\,K^{\mathrm{cone}}_{j}\,\Phi^{<\rho}\otimes\Psi^{<\rho}\,d\mu^{\otimes 2}_{g^{\mathrm{cone}}_{j}}\Bigg].$ (2)
@@
1.5 Main result: renormalized bilinear limit across the discriminant
@@
$\lim_{t\downarrow 0}\ \iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau),$ (3)
@@ 1.6 Idea of the proof
1.7 Further consequences and organization of the paper
@@
2 Semi-flat collapse and normalized identifications
@@
2.1 Geometric setting on the regular locus
@@
2.2 Semi-flat product control and collapse scale
@@
$\mathrm{diam}_{g_{F,t}}(F_{b,t})\leq s(t)\qquad\text{for all }b\in B_{0}.$ (5)
@@
2.3 Quantitative disintegration of volume
@@
$\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$ (7)
@@
$\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$ (8)
@@
$\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}\ \leq\ C\,\varepsilon_{t}.$ (10)
@@
2.4 Normalized lift and fiber-average operators
@@
For adjointness, let $\phi\in L^{2}(B_{0})$ and $u\in L^{2}(X_{t,0})$. Using (7),
$\langle I_{t}\phi,u\rangle_{L^{2}(X_{t,0})}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}A_{t}(b)^{-1/2}\phi(b)\,u\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b)=\int_{B_{0}}\phi(b)\,(P_{t}u)(b)\,d\mu_{B}(b)=\langle\phi,P_{t}u\rangle_{L^{2}(B_{0})}.$
@@
2.5 Horizontal $H^{1}$ compatibility for fiber-constant lifts
@@
$\int_{X_{t,0}}|\nabla I_{t}v|_{g(t)}^{2}\,d\mu_{t}=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B}\ +\ O(\varepsilon_{t})\Big(\|v\|_{H^{1}(B_{0})}^{2}\Big).$ (16)
@@
2.6 Fiber-constant projection (notation)
@@
2.7 Section summary and next steps
@@
3 Dirichlet forms, vertical Poincaré inequality, and Mosco convergence
@@
3.1 Dirichlet forms on the base and on the total space
@@
$E_{B}[v]:=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B},\qquad\mathrm{Dom}(E_{B})=H^{1}_{0}(B_{0}).$ (17)
@@
$E_{t}[u]:=\int_{X_{t,0}}|\nabla u|_{g(t)}^{2}\,d\mu_{t},\qquad\mathrm{Dom}(E_{t})=H^{1}_{0}(X_{t,0}).$ (18)
@@
3.2 Normalized identifications and the fiber-constant projection
@@
$\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$ (19)
@@
$\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$ (20)
@@
$\|\rho_{t}-1\|_{L^{\infty}(X_{t,0})}+\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}+\|\nabla_{B}\rho_{t}\|_{L^{\infty}(X_{t,0})}\ \leq\ C\,\varepsilon_{t},$ (21)
@@
3.3 Vertical Poincaré inequality from the spectral gap
@@
$\int_{F_{b,t}}\Big|u-\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}\Big|^{2}\rho_{t}\,d\vartheta_{b,t}\ \leq\ C\,s(t)^{2}\int_{F_{b,t}}|\nabla_{F}u|_{g_{F,t}}^{2}\,\rho_{t}\,d\vartheta_{b,t}.$ (25)
@@
$\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}\ \leq\ C\,s(t)^{2}\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}.$ (26)
@@
$\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}\ \geq\ c\,s(t)^{-2}\,\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}.$ (27)
@@
3.4 Energy comparison for lift and average
@@
3.5 Mosco convergence of Dirichlet forms on identified spaces
@@
$\liminf_{t\downarrow 0}E_{B}[v_{t}]\ \leq\ \liminf_{t\downarrow 0}E_{t}[u_{t}].$ (31)
@@ 3.6 Section summary
4 From Mosco convergence to compressed semigroups and the interior bilinear kernel limit
@@
[…]
@@
4.1 Generators and the compressed operators
@@
$\|S_{t}(\tau)\|_{L^{2}_{B}\to L^{2}_{B}}\leq 1,\qquad S_{t}(\tau)=S_{t}(\tau)^{*}.$ (33)
4.2 Mosco convergence implies strong resolvent convergence
@@
$P_{t}(\alpha+H_{t})^{-1}I_{t}f\longrightarrow(\alpha+H_{B})^{-1}f\qquad\text{strongly in }L^{2}_{B}.$ (34)
@@
4.3 Strong convergence of compressed heat operators
@@
$S_{t}(\tau)=P_{t}e^{-\tau H_{t}}I_{t}\longrightarrow e^{-\tau H_{B}}=T_{B}(\tau)\qquad\text{strongly in }\mathcal{B}(L^{2}_{B}).$ (35)
@@
4.4 Heat kernels and bilinearization
@@
$\langle F,T_{t}(\tau)G\rangle_{L^{2}_{t}}$ $=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,F(x)\,G(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$ (36)
$\langle\phi,T_{B}(\tau)\psi\rangle_{L^{2}_{B}}$ $=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$ (37)
@@
$\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\langle I_{t}\Phi,T_{t}(\tau)I_{t}\Psi\rangle_{L^{2}_{t}}.$ (38)
@@
$\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$ (39)
@@
$\langle\Phi,T_{B}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$ (40)
@@
4.5 Interior bilinear heat-kernel pairing limit

### v2

# Heat Flow under Semi-Flat Collapse with Conic Renormalization

@@
Abstract.
Motivated by the Strominger–Yau–Zaslow (SYZ) collapse of elliptic K3 surfaces, we study heat flow on Ricci-flat Kähler manifolds $(X_{t},g(t))$ equipped over $B_{\mathrm{reg}}=B\setminus D$ with a semi-flat $2$-torus fibration $\Pi_{t}:X_{t}\to B$. Assuming an exponentially accurate semi-flat product approximation and a uniform vertical spectral gap at the collapse scale, we prove that for each fixed $\tau>0$ the fiber-compressed heat operators $P_{t}e^{-\tau\Delta_{g(t)}}I_{t}$ converge strongly on $L^{2}$ to the base heat semigroup $e^{-\tau\Delta_{g_{B}}}$; equivalently, bilinear pairings of the total-space heat kernel against fiber-constant test functions converge on compact subsets of $B_{\mathrm{reg}}$.
Near the discriminant $D$ the limiting base metric is conic. We introduce a canonical renormalized bilinear functional $K_{B}^{\mathrm{ren}}$ by replacing the pairing in neighborhoods of each conic point with the model flat-cone heat kernel. A conic parametrix proves cutoff- and chart-independence of $K_{B}^{\mathrm{ren}}$ and shows that the total-space pairings converge to $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ for all $\Phi,\Psi\in C_{c}^{\infty}(B)$. Under exponentially small geometric errors we obtain an explicit exponential convergence rate, uniform for $\tau$ in compact subsets of $(0,\infty)$, and we also treat Neumann and nonnegative Robin boundary conditions.
1. Introduction and statement of results
1.1. Motivation: heat flow under SYZ-type collapse
@@
1.2. Geometric setting and analytic identifications
@@
[…]
Figure 1. Schematic picture of the fibration $\Pi_{t}:X_{t}\to B$ with finite discriminant set $D$. Over the regular locus $B_{\mathrm{reg}}=B\setminus D$, the fibers are flat $2$-tori $\mathbb{T}^{2}$ that collapse as $t\downarrow 0$.
@@
1.3. Bilinear pairings and compressed semigroups
@@
(1) $\iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ \langle\Phi,\ (P_{t}T_{t}(\tau)I_{t})\Psi\rangle_{L^{2}(B_{0})}.$
@@
1.4. Discriminant regime and intrinsic conic renormalization
@@
(2) $\begin{split}K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[&\iint_{B\times B}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\Psi^{(\rho)}(b^{\prime})\\ &\qquad d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})\\ &+\sum_{j=1}^{N}\iint_{U_{j}\times U_{j}}(\chi_{j}\otimes\chi_{j})\,K^{\mathrm{cone}}_{j}\\ &\qquad\Phi^{<\rho}\otimes\Psi^{<\rho}\,d\mu^{\otimes 2}_{g^{\mathrm{cone}}_{j}}\Bigg].\end{split}$
@@
1.5. Main result: renormalized bilinear limit across the discriminant
@@
(3) $\lim_{t\downarrow 0}\ \iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau),$
@@ 1.6. Idea of the proof
1.7. Further consequences and organization of the paper
@@
2. Semi-flat collapse and normalized identifications
@@
2.1. Geometric setting on the regular locus
@@
2.2. Semi-flat product control and collapse scale
@@
(5) $\mathrm{diam}_{g_{F,t}}(F_{b,t})\leq s(t)\qquad\text{for all }b\in B_{0}.$
@@
2.3. Quantitative disintegration of volume
@@
(7) $\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$
@@
(8) $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$
@@
(10) $\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}\ \leq\ C\,\varepsilon_{t}.$
@@
2.4. Normalized lift and fiber-average operators
@@
For adjointness, let $\phi\in L^{2}(B_{0})$ and $u\in L^{2}(X_{t,0})$. Using (7), we compute
$\langle I_{t}\phi,u\rangle_{L^{2}(X_{t,0})}$ $=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}A_{t}(b)^{-1/2}\phi(b)\,u\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b)$
$=\int_{B_{0}}\phi(b)\,(P_{t}u)(b)\,d\mu_{B}(b)$
(14) $=\langle\phi,P_{t}u\rangle_{L^{2}(B_{0})}.$
@@
2.5. Horizontal $H^{1}$ compatibility for fiber-constant lifts
@@
(17) $\int_{X_{t,0}}|\nabla I_{t}v|_{g(t)}^{2}\,d\mu_{t}=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B}\ +\ O(\varepsilon_{t})\Big(\|v\|_{H^{1}(B_{0})}^{2}\Big).$
@@
2.6. Fiber-constant projection (notation)
@@
2.7. Section summary and next steps
@@
3. Dirichlet forms, vertical Poincaré inequality, and Mosco convergence
@@
3.1. Dirichlet forms on the base and on the total space
@@
(18) $E_{B}[v]:=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B},\qquad\mathrm{Dom}(E_{B})=H^{1}_{0}(B_{0}).$
@@
(19) $E_{t}[u]:=\int_{X_{t,0}}|\nabla u|_{g(t)}^{2}\,d\mu_{t},\qquad\mathrm{Dom}(E_{t})=H^{1}_{0}(X_{t,0}).$
@@
3.2. Normalized identifications and the fiber-constant projection
@@
(20) $\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$
@@
(21) $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$
@@
(22) $\|\rho_{t}-1\|_{L^{\infty}(X_{t,0})}+\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}+\|\nabla_{B}\rho_{t}\|_{L^{\infty}(X_{t,0})}\ \leq\ C\,\varepsilon_{t},$
@@
3.3. Vertical Poincaré inequality from the spectral gap
@@
$\Longrightarrow\text{suppression of nonconstant fiber modes at low energy}.$
@@
(26) $\int_{F_{b,t}}\Big|u-\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}\Big|^{2}\rho_{t}\,d\vartheta_{b,t}\ \leq\ C\,s(t)^{2}\int_{F_{b,t}}|\nabla_{F}u|_{g_{F,t}}^{2}\,\rho_{t}\,d\vartheta_{b,t}.$
@@
(27) $\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}\ \leq\ C\,s(t)^{2}\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}.$
@@
(28) $\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}\ \geq\ c\,s(t)^{-2}\,\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}.$
@@
3.4. Energy comparison for lift and average
@@
3.5. Mosco convergence of Dirichlet forms on identified spaces
@@
(32) $\liminf_{t\downarrow 0}E_{B}[v_{t}]\ \leq\ \liminf_{t\downarrow 0}E_{t}[u_{t}].$
@@ 3.6. Section summary
4. From Mosco convergence to compressed semigroups and the interior bilinear kernel limit
@@
$\text{Mosco convergence of }E_{t}$ $\Longrightarrow\text{strong resolvent/semigroup convergence}$
$\Longrightarrow\text{bilinear heat-kernel pairing limit}$
$\qquad\text{on }B_{0}\Subset B_{\mathrm{reg}}.$
@@
4.1. Generators and the compressed operators
@@
(34) $\|S_{t}(\tau)\|_{L^{2}_{B}\to L^{2}_{B}}\leq 1,\qquad S_{t}(\tau)=S_{t}(\tau)^{*}.$
4.2. Mosco convergence implies strong resolvent convergence
@@
(35) $P_{t}(\alpha+H_{t})^{-1}I_{t}f\longrightarrow(\alpha+H_{B})^{-1}f\qquad\text{strongly in }L^{2}_{B}.$
@@
4.3. Strong convergence of compressed heat operators
@@
(36) $S_{t}(\tau)=P_{t}e^{-\tau H_{t}}I_{t}\longrightarrow e^{-\tau H_{B}}=T_{B}(\tau)\qquad\text{strongly in }\mathcal{B}(L^{2}_{B}).$
@@
4.4. Heat kernels and bilinearization
@@
(37) $\langle F,T_{t}(\tau)G\rangle_{L^{2}_{t}}$ $=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,F(x)\,G(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$
(38) $\langle\phi,T_{B}(\tau)\psi\rangle_{L^{2}_{B}}$ $=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
@@
(39) $\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\langle I_{t}\Phi,T_{t}(\tau)I_{t}\Psi\rangle_{L^{2}_{t}}.$
@@
(40) $\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$
@@
(41) $\langle\Phi,T_{B}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
@@
4.5. Interior bilinear heat-kernel pairing limit